OpenStudy (anonymous):

the volume of a frustrum of a right circular cone is 52pi ft^3. Its altitude is 3 ft and measure of its lower radius is three times the measure of its upper radius. Find the lateral area of the frustum.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

\[1/3 * \Pi h(R^2 + r^2 + Rr)\] Volume of frustrum is given by (formula drawn below) where R and r are the lower and upper base radius.

OpenStudy (anonymous):

what can we do to get the values of R and r?

OpenStudy (anonymous):

@perl

OpenStudy (igreen):

\(\sf V = \dfrac{\pi h}{3} (R^2 + Rr + r^2)\) 'h' is the height, 'R' is the radius of the lower base, and 'r' is the radius of the upper base.

OpenStudy (igreen):

Plug in what we know: \(\sf 52\pi = \dfrac{\pi (3)}{3} (R^2 + Rr + r^2)\) We can cancel out pi. \(\sf 52 = \dfrac{3}{3} (R^2 + Rr + r^2)\) Simplify: \(\sf 52 = (R^2 + Rr + r^2)\) Okay, it says the lower radius is 3 times the measure of the upper radius. So: R = 3r

OpenStudy (igreen):

Plug in 3r for 'R' and solve for 'r'

OpenStudy (igreen):

\(\sf 52 = ((3r)^2 + (3r)r + r^2)\)

OpenStudy (igreen):

Can you simplify that? @~Gelmhar

OpenStudy (anonymous):

r=2.73

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

\[7r^2=52 , r^2=52/7 ,r=2.73\]

OpenStudy (igreen):

No.. I got something else.

OpenStudy (anonymous):

how?

OpenStudy (igreen):

\(\sf 54 = ((3r)^2 + (3r)r + r^2)\) \(\sf (3r)^2 \rightarrow 3^2r^2 \rightarrow 9r^2\) \(\sf (3r)r \rightarrow 3r \times r \rightarrow 3r^2\) So we have: \(\sf 54 = (9r^2 + 3r^2 + r^2)\) Combine like terms.

OpenStudy (anonymous):

i see.. wait i will solve it

OpenStudy (anonymous):

r=2

OpenStudy (anonymous):

can you help me again @iGreen

OpenStudy (igreen):

Yep, you got it.

OpenStudy (igreen):

\(54=(9r^2+3r^2+r^2)\) \(54 = 13r^2\) \(4 = r^2\) \(r = 2\)

OpenStudy (igreen):

\(\sf LA=\pi(R+r)L\)

OpenStudy (igreen):

@~Gelmhar Do you have a picture?

OpenStudy (anonymous):

i got the answer now, lets move to another problem can we ? :D

OpenStudy (igreen):

Great